A 200 g set of keys , sitting at rest on a table falls 175 cm to the ground. Ignoring frictional forces. How much is the initial kinetic energy of the keys ? How much is the initial potential energy ? How much is the final kinetic energy ? How much is the final potential energy ? How much is the velocity right before it hits the ground ?


1
Expert's answer
2021-01-31T14:03:36-0500

a) Initial kinetic energy of the keys equals zero.


"KE_i=0."


b) Let's calculate the initial potential energy of the keys:


"PE_i=mgh=0.2\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot 1.75\ m=3.43\ J."

c) According to the law of conservation of energy final kinetic energy equals initial potential energy:


"KE_f=PE_i=3.43\ J."

d) Final potential energy equals zero:


"PE_f=0."

e)

"PE_i=KE_f,""mgh=\dfrac{1}{2}mv^2,""v=\sqrt{2gh}=\sqrt{2\cdot9.8\ \dfrac{m}{s^2}\cdot1.75\ m}=5.86\ \dfrac{m}{s}."